What is the equation in slope intercept form of the line that goes through the coordinates of (-4,6) (8,12)?
Answer by Edwin McCravy(20086)

What is the equation in slope intercept
form of the line that goes through the coordinates of 
(-4,6) (8,12)?

Use the slope formula:

     y2 - y1
m = —————————
     x2 - x1

where (x1, y1) = (-4, 6) and (x2, y2) = (8, 12)

     (12) - (6)      6      6
m = ———————————— = ————— = ———— = 1/2
     (8) - (-4)     8+4     12 

Now substitute in the point slope formula:

   y - y1 = m(x - x1)

   y - 6 = (1/2)[x - (-4)]

Multiply both sides by LCD = 2

2(y - 6) = 2(1/2)(x + 4)

 2y - 12 = 1(x + 4)

 2y - 12 = x + 4

      2y = x + 16

Multiply thru by 1/2

       y = (1/2)x + 8       

Edwin